OpenStudy (anonymous):

Show that the set \[\left\{ 1,x,x^2,x^3,... \right\}\]is a basis of the vector space \[\mathbb{R}[x]\]

OpenStudy (anonymous):

\(\mathbb{R}[x]\) is the vector space of n'th degree polynomials involving one variable?

OpenStudy (anonymous):

Just not sure about the notation I'm afraid..

OpenStudy (anonymous):

im trying to think of how to word this lol <.<

OpenStudy (anonymous):

um... ok. you want to show that if you have a linear combination of the basis "vectors", that: \[c_0(1)+c_1(x)+c_2(x^2)+\ldots c_n(x^n) = 0 \iff c_0 =c_1 = \ldots = c_n = 0 \] Then you claim that this is true for any n.

OpenStudy (anonymous):

now that first statement: \[c_0(1)+c_1(x)+c_2(x^2)+…+c_n(x^n)=0\] has to hold for all x. By the Fundamental Theorem of Algebra, we know that a polynomial of degree n has at most n roots. So for the above to hold for all x, that does imply that each coefficient is 0. So the set: \[\left\{ 1,x,x^2,\ldots,x^n \right\}\]is linearly independent, and because this holds for all n, the set: \[\left\{ 1,x,x^2,\ldots \right\}\]is also linearly independent.

OpenStudy (anonymous):

You also have to satisfy the other requirements of a basis.

OpenStudy (anonymous):

Closure under vector addition, and closure under scalar multiplication.

OpenStudy (anonymous):

those are the requirements for a space to be a vector space. I dont think that has anything to do with the basis itself. \[\mathbb R[x]\] does follow those 2 properties (along with 0 being in the space as well.) It is the set of all polynomials in one variable x, whose coefficients are real numbers.